Consider the following configuration of solar photovoltaic arrays consisting of crystalline silicon solar cells There are two subsystems connected in parallel, each one containing two cells. In order for the system to function, at least one of the two Within each subsystem, the two cells are connected in series, so a subsystem will work only if all cells in the subsystem wor suppose i since the cells are identical. parallel subsystems must work he lifetime of cell / exceeds t, 0 1,2. ). We assume we want to determine the probability that the system lifetime exceeds t Let A, denote the event that t that the As are Independent events (whether any particular cell lasts more than to hours has no bearing on to hours has no bearing on whether or not any other cell does) and that PLA) - 0.6 for every - 0.6, the probability that system lifetime exceeds to is easily seen to be 0.5904. To what value would 0.6 have to be changed in order to increase the systenm retim o rela tity from o ss04 to o 61, н ti let p A)-р, express system rea tility in terms of p, and then let x-pA 0 3485 (Round your answer to four decir al places.) Need Help? Tat
Answer #1

here as from above probability that system lifetime increase from to =1-(1-p2)2 =0.61

(1-p2)2 =0.39

solving above p=0.6128